OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <6, 4>, v = <7, 5>

OpenStudy (anonymous):

familiar with dot product theorem?

OpenStudy (zzr0ck3r):

\(<6,4>\cdot<7,5>= |u||v|cos(\theta)\) so \(6*7+4*5 = |u||v|cos(\theta)\) so \(\frac{62}{|u||v|}=\cos(\theta)\)

OpenStudy (anonymous):

would the answer be 11.8 degrees?

OpenStudy (zzr0ck3r):

\(|u|=\sqrt{6^2+4^2}=\sqrt{44}\) \(|v|=\sqrt{7^2+5^2}=\sqrt{74}\)

OpenStudy (zzr0ck3r):

\(\cos(\theta)=\frac{62}{\sqrt{44}\sqrt{74}}\) so \(\theta=\cos^{-1}(\frac{62}{\sqrt{44*74}}) \approx0.413\)

OpenStudy (anonymous):

That's not one of my answer choices. 1.8° -9.1° 0.9° 11.8°

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

woops \(\sqrt{6^2+4^2}=\sqrt{52}\)

OpenStudy (anonymous):

so 1.8 degrees

OpenStudy (zzr0ck3r):

\(\theta=\cos^{-1}(\frac{62}{\sqrt{52*74}}) \approx 0.0322\approx1.8^{\circ} \)

OpenStudy (anonymous):

wonderful, thanks!

OpenStudy (zzr0ck3r):

np